# Statistical Quantum Mechanics of the Random Permutation Sorting System (RPSS): A Self-Stabilizing True Uniform RNG

### v1

@@ Abstract
We introduce the Random Permutation Sorting System (RPSS), a novel, software-defined framework that generates high-quality, uniform randomness by leveraging a statistical model with properties analogous to those of quantum ensembles. The RPSS operates as a self-contained entropy engine based on a fundamental pair of conjugate observables: the permutation count ($\hat{N}_{p}$) and the elapsed sorting time ($\hat{T}$). While the raw distributions of these observables are highly right-skewed and heavy-tailed, we provide a mathematical proof of their synchronous convergence to uniformity. This convergence arises from the compounded effect of permutation counts on elapsed time and a process of modular uniformization, in which a vast number of microstates—arising from combinatorial complexity and intrinsic system-level jitter—map onto a finite set of uniform symbols. This mechanism enables the RPSS to function as a True Uniform Random Number Generator (TURNG). We implemented the RPSS as the QPP-RNG, a practical software realization that autonomously harvests entropy from a platform’s intrinsic microarchitectural dynamics. Unlike prior benchmark-focused studies, this work provides a deep theoretical and experimental investigation of the underlying physical mechanism. Empirical validation confirms our predictions, showing rapid entropy convergence and statistically uniform outputs, as assessed by NIST SP 800-90B min-entropy and chi-squared statistics, consistent with central-limit-theorem expectations, as demonstrated in our previous publications. RPSS establishes a new class of quantum-inspired eco-cryptosystems, in which randomness is simultaneously harvested from unpredictable hardware behavior and amplified through a combinatorial process. This dual-source approach provides a compact, self-stabilizing entropy engine suitable for foundational applications in cryptography, blockchain protocols, and digital currency systems, offering a robust, platform-agnostic alternative to conventional entropy sources.
@@ I Introduction
Building on the framework of the Quantum Permutation Pad (QPP) [15, 16], which generalizes the one-time pad into the full permutation group of $n$-bit data, we recently developed QPP-RNG: a software-based TRNG harvesting entropy from microarchitectural fluctuations in commodity hardware [17, 18]. These fluctuations, measured in discrete system clock units (ticks or nanoseconds), arise from CPU pipeline jitter, cache misses, memory-access latency, and OS scheduling. By cycling through iterative permutations, QPP-RNG amplifies this entropy and produces statistically uniform outputs independently of external post-processing.
@@ II.3.1 Permutation Count
Let $M=m\cdot N!$ represent the effective size of the permutation space when considering $m$ independent permutation groups. Random sorting over $m$ permutation groups is equivalent to sorting a single permutation group with success probability $p=1/M$.
@@
with success probability $p=1/M=1/(m\cdot N!)$.
@@
$\mathbb{E}[n_{p}]=\frac{m}{p}=M,\qquad\mathrm{Var}(n_{p})=\frac{m(1-p)}{p^{2}}\approx M^{2}.$
@@
$\mathbb{E}[n_{p}]=\frac{1}{p}=M,\qquad\mathrm{Var}(n_{p})=\frac{1-p}{p^{2}}\approx M^{2}.$
In the continuous limit with $M\gg 1$ and $p\ll 1$, the geometric law converges to an exponential distribution with rate parameter $\lambda=1/M$:
$\Pr[n_{p}\approx x]\sim\lambda e^{-\lambda x},\qquad\lambda=\frac{1}{M}.$ (14)
This theoretical result agrees with empirical observations (e.g., for $N>2$, $m=1$), where the histogram of $n_{p}$ decays exponentially, with $\Pr[n_{p}=1]\approx 1/M$.
@@
  • For $m=1$, most realizations of $n_{p}$ cluster around $M=N!$, with the probability mass decaying exponentially.
  • For $m>1$, the distribution broadens: realizations cluster around $M=mN!$, while the probability mass decays more slowly for large values of $n_{p}$. This matches empirical histograms where the exponential-like shape of the $m=1$ case evolves into a longer-tailed distribution as $m$ increases.
The unified parameter $M=m\cdot N!$ serves as the characteristic scale of the distribution, determining both its mean and variance. This formulation emphasizes the continuous transition between the geometric ($m=1$) and negative binomial ($m>1$) cases, providing a consistent framework for analyzing the permutation count statistics in the RPSS system.
@@ II.3.2 Elapsed-Time Observable $\hat{T}$
As $m$ increases, the distribution of $\hat{N}_{p}$ broadens. Consequently, the distribution of $\hat{T}$ evolves from the sharp, discrete-like shape of the $m=1$ case into a smooth, right-skewed distribution. For $m=2,3,4$, $\hat{T}$ exhibits a characteristic right-skew with a peak that shifts right to approximately $m\cdot\mathbb{E}[X]$ and a long tail that decays more slowly than that of $\hat{N}_{p}$ alone, matching the empirical histograms. The convolution with the runtime distribution $X_{j}$ acts to smooth the distribution and extend its tail.
@@ III Experiments and Discussions
Experiments were performed on the following hardware and software platform:
Table 1: Experimental platform: a RPSSnsystem
@@ III.3 Modular-Reduced Distributions of the Conjugate Observables
Figure 3: Modular-reduced permutation count distributions for sorting the disordered array $\{3,2,0,1\}$ at repetition values $m=1,2,3$.
@@
Figure 4: Modular-reduced elapsed-time distributions (in ticks) for sorting the disordered array $\{3,2,0,1\}$ at repetition values $m=1,2,3$.
@@
3 15 90 $n=4$ Good
3 20 120 $n=4$ Excellent
4 3 72 $n=4$ Good
4 4 96 $n=4$ Excellent
5 2 240 $n=4$ Excellent
5 4 480 $n=8$ Good
5 5 600 $n=8$ Excellent
@@ IV Conclusion
The empirical validation confirms our theoretical predictions. The compound nature of the elapsed-time observable, which incorporates real-world system noise, is a powerful source of unpredictability. The modular uniformization process, in which a large number of microstates collapse onto a small set of computational symbols, acts as a cryptographic extractor, eliminating residual correlations and producing outputs that pass stringent statistical tests. Our analysis of entropy convergence as a function of array size ($N$) and repetition parameter ($m$) provides a practical guide for implementing a self-stabilizing entropy source on commodity hardware.

### v2

@@ Abstract
We introduce the Random Permutation Sorting System (RPSS), a novel, software-defined framework that generates high-quality, uniform randomness by leveraging a statistical model with properties analogous to quantum ensembles. The RPSS operates as a self-contained entropy engine based on a fundamental pair of conjugate observables: the permutation count ($\hat{N}_{p}$) and the elapsed sorting time ($\hat{T}$). While the raw distributions of these observables are highly right-skewed and heavy-tailed, we provide a mathematical proof of their synchronous convergence to uniformity. The permutation count distributions follow a negative binomial model, and the observed modes for different repetition parameters $m$ closely match the theoretical predictions, confirming the validity of the model. This convergence arises from the compounded effect of permutation counts on elapsed time and a process of modular uniformization, in which a vast number of microstates—stemming from combinatorial complexity and intrinsic system-level jitter—map onto a finite set of uniform symbols. This mechanism enables the RPSS to function as a True Uniform Random Number Generator (TURNG). We implemented the RPSS as the QPP-RNG, a practical software realization that autonomously harvests entropy from a platform’s intrinsic microarchitectural dynamics. Unlike prior benchmark-focused studies, this work provides a deep theoretical and experimental investigation of the underlying physical mechanism. Empirical validation confirms our predictions, demonstrating rapid entropy convergence and statistically uniform outputs, as assessed by NIST SP 800-90B min-entropy and chi-squared statistics, consistent with central-limit-theorem expectations. The RPSS establishes a new class of quantum-inspired eco-cryptosystems, in which randomness is simultaneously harvested from unpredictable hardware behavior and amplified through a combinatorial process. This dual-source approach provides a compact, self-stabilizing entropy engine suitable for foundational applications in cryptography, blockchain protocols, and digital currency systems, offering a robust, platform-agnostic alternative to conventional entropy sources.
@@ I Introduction
Building on the framework of the Quantum Permutation Pad (QPP) [5, 10], which generalizes the one-time pad into the full permutation group of $n$-bit data, we recently developed QPP-RNG: a software-based TRNG harvesting entropy from microarchitectural fluctuations in commodity hardware [26]. These fluctuations, measured in discrete system clock units (ticks or nanoseconds), arise from CPU pipeline jitter, cache misses, memory-access latency, and OS scheduling. By cycling through iterative permutations, QPP-RNG amplifies this entropy and produces statistically uniform outputs independently of external post-processing.
@@
To provide a clear and quantitative basis for describing the observed uniformity, we classify QPP-RNG outputs using the NIST SP 800-90B MCV (Most Common Value) min-entropy estimator.
Streams with min-entropy above 7.95 bits/byte are labeled as Near uniform, and those above 7.99 bits/byte as Essential uniform. Supporting statistics such as mean and serial correlation are reported for diagnostic purposes but do not determine classification (see Sec. III).
@@ II.3.1 Permutation Count
Let $M=m\cdot N!$ represent the effective counting for the $m$-th success. However, the success probability per trial is determined solely by the size of the permutation group:
$p=\frac{1}{N!}.$
This reflects the probability that a single permutation trial produces the inverse permutation $\hat{p}^{-1}$.
@@
with success probability $p=1/N!$.
@@
$\mathbb{E}[n_{p}]=\frac{m}{p}=m\cdot N!,\qquad\mathrm{Var}(n_{p})=\frac{m(1-p)}{p^{2}}\approx m\cdot(N!)^{2}.$
@@
$\mathbb{E}[n_{p}]=\frac{1}{p}=N!,\qquad\mathrm{Var}(n_{p})=\frac{1-p}{p^{2}}\approx(N!)^{2}.$
In the continuous limit with $N!\gg 1$ and $p\ll 1$, the geometric law converges to an exponential distribution with rate parameter $\lambda=1/N!$:
$\Pr[n_{p}\approx x]\sim\lambda e^{-\lambda x},\qquad\lambda=\frac{1}{N!}.$ (14)
This theoretical result aligns with empirical observations (e.g., for $N>2$, $m=1$), where the histogram of $n_{p}$ decays approximately exponentially, with the first trial probability $\Pr[n_{p}=1]=p=1/N!$.
Mode of Permutation Count Distribution
In addition to the expectation and variance, we also examine the mode of the permutation count distribution. The mode corresponds to the value of $n_{p}$ that occurs most frequently, i.e., the most likely number of trials required for the system to generate the inverse permutation $\hat{p}^{-1}$.
For the general case of the $m$-th success ($m>1$), the mode of the negative binomial distribution is given by
$n_{p}^{\text{mode}}=\left\lfloor\frac{(m-1)(1-p)}{p}\right\rfloor+1,$
where $p$ is the success probability per trial. In our RPSS system, each permutation has a probability of
$p=\frac{1}{N!}.$
Substituting this into the mode formula, we obtain
$n_{p}^{\text{mode}}=(m-1)(N!-1)+1.$ (15)
This accurately predicts the most likely number of trials required for the $m$-th occurrence of the inverse permutation.
For the special case of $m=1$ (first success), the distribution reduces to a geometric distribution, whose mode is
$n_{p}^{\text{mode}}=1,$
indicating that the most frequent number of trials is just the first trial. This aligns with the observed RPSS behavior, where for $N=4$ the modes for $m=2,3,4$ are observed as 24–25, 48–49, and so on, consistent with the theoretical predictions.
@@
  • For $m=1$, most realizations of $n_{p}$ cluster around $M=N!$, with the probability mass decaying exponentially. The mode of the distribution is also approximately $M$, reinforcing this behavior.
  • For $m>1$, the distribution broadens: realizations cluster around $M=mN!$, while the probability mass decays more slowly for large values of $n_{p}$. The mode shifts to (m1)M(m-1)M, and the distribution becomes longer-tailed.
The unified parameter $M=m\cdot N!$ serves as the characteristic scale of the distribution, determining both its mean, variance, and mode. This formulation emphasizes the continuous transition between the geometric ($m=1$) and negative binomial ($m>1$) cases, providing a consistent framework for analyzing the permutation count statistics in the RPSS system.
@@ II.3.2 Elapsed-Time Observable $\hat{T}$
As $m$ increases, the distribution of $\hat{N}_{p}$ broadens. Consequently, the distribution of $\hat{T}$ evolves from the sharp, discrete-like shape observed for $m=1$ into a smooth, right-skewed distribution. For $m=2,3,4$, $\hat{T}$ exhibits a characteristic right skew, with a peak that shifts right in proportion to $m$, consistent with the scaling of $\mathbb{E}[\hat{T}$]. The convolution with the runtime distribution $X_{j}$ smooths the distribution and extends its right tail, in agreement with the empirical histograms.
@@ III Experiments and Discussions
Experiments were performed on the following hardware and software platform summarized in Table 1.
Table 1: Experimental platform: an RPSS system
@@ III.2 Raw Distributions of the Conjugate Observables
In our experiments for $N=4$, the observed modes of the permutation count distributions were 1 for $m=1$, 22 for $m=2$, 48 for $m=3$, and 71 for $m=4$. The theoretical mode for $m>1$ is given by Eq. (15). Table 4 compares the observed and theoretical modes for $m=1,2,3,4$:
Table 4: Comparison of observed and theoretical modes of the permutation count distribution for $N=4$.
$m$ Observed Mode Theoretical Mode (Eq. (15))
1 1 1
2 22 24
3 48 47
4 71 70
The small discrepancies for $m=2$ and $m=3$ are due to the right-skewness of the distribution and the discrete nature of $n_{p}$, which slightly shifts the empirical peak relative to the theoretical value. Overall, these results confirm that the negative binomial framework accurately models the RPSS permutation count distribution, capturing the progressive rightward shift of the mode and the elongation of the tail as $m$ increases.
@@ III.3 Modular-Reduced Distributions of the Conjugate Observables
Figure 3: Modular-reduced permutation count distributions for sorting the disordered array $\{3,2,0,1\}$ at repetition value $m=3$.
@@
Figure 4: Modular-reduced elapsed-time distributions (in ticks) for sorting the disordered array $\{3,2,0,1\}$ at repetition value $m=3$.
@@
3 15 90 $n=4$ Near
3 20 120 $n=4$ Essential
4 3 72 $n=4$ Near
4 4 96 $n=4$ Essential
5 2 240 $n=4$ Essential
5 4 480 $n=8$ Near
5 5 600 $n=8$ Essential
Clarification of Near-Uniform and Essentially Uniform Behavior
To remove ambiguity in terminology, we classify QPP-RNG outputs using the NIST SP 800-90B MCV min-entropy estimator. Since min-entropy lower-bounds Shannon entropy, it provides a conservative and reliable measure of uniformity. A stream is considered essentially uniform if its estimated min-entropy exceeds 7.99 bits/byte, and near-uniform if it exceeds 7.95 bits/byte. Supporting statistics, such as the arithmetic mean and serial correlation, are reported for diagnostic purposes but are not used for classification.
Table 6: Min-entropy-based uniformity metrics for the 4-element QPP-RNG array.
$m$ $N_{s}$ Elapsed Time Permutation Uniformity Class
4 $10^{6}$ 7.932 7.931 Near
4 $10^{7}$ 7.977 7.981 Essential
4 $10^{8}$ 7.994 7.994 Essential
3 $10^{6}$ 7.905 7.933 Near
3 $10^{7}$ 7.973 7.975 Essential
3 $10^{8}$ 7.979 7.990 Essential
Table 6 demonstrates that QPP-RNG outputs transition predictably from near-uniform to essentially uniform behavior as the permutation depth $m$ or sample size $N_{s}$ increases, providing a quantitative and reproducible framework for assessing uniformity.
@@ III.4 QPP-RNG as a Self-Training True Uniform RNG (TURNG)
III.5 Comparison with Representative RNG Methods
Table 7 places RPSS in context with representative TRNG and QRNG implementations whose min-entropy has been evaluated using comparable NIST SP 800-90B methodologies. While many modern RNGs successfully pass standard statistical test suites, these results illustrate that test pass rates alone do not fully capture architectural differences in entropy generation mechanisms.
Table 7: Comparative Min-Entropy $\mathbf{H_{\min}}$ Results
RNGs Class bits/bit bits/byte
RPSS (this work) TURNG 0.9974 7.9792
Red Hat RNG [20] CPU Jitter 0.9316 7.4528
Microchip [19] TRNG 0.5071 4.0568
van der Waals [1] TRNG 0.9830 7.8640
Tunneling Effect [12] QRNG 0.9872 7.8976
IDQ Quantis [18] QRNG 0.9843 7.8744
Quside PCIe [21] QRNG 0.8142 6.5136
We emphasize that the reported min-entropy values for validated CPU jitter RNGs (e.g., Red Hat RNG [20]) correspond to entropy estimates after conservative conditioning or entropy crediting at the module boundary, as required by NIST SP 800-90B validation. These values therefore reflect effective post-conditioning entropy rather than intrinsic raw-source entropy. In contrast, the RPSS/QPP-RNG results reported in Table 7 correspond to the native output of the generator and do not rely on external conditioning, extractors, or whitening stages.
Conventional CPU jitter–based RNGs rely on stochastic microarchitectural effects and typically require explicit conditioning or conservative entropy crediting at the module interface, since raw jitter measurements alone do not guarantee uniform symbol distributions. Physical TRNGs based on electronic or material noise sources (e.g., Microchip ECC608 or van der Waals heterostructures) demonstrate strong entropy generation but depend on specialized hardware and analog components. QRNGs achieve high min-entropy by exploiting intrinsic quantum indeterminacy, though at the cost of optical or solid-state quantum hardware and associated system complexity.
In contrast, RPSS achieves Near to Essential uniformity through an internal structural mechanism: entropy harvested from timing jitter is amplified by permutation degeneracy and modular projection, without reliance on external extractors or post-processing. This distinguishes RPSS from both software jitter RNGs and hardware-based TRNG/QRNG systems. As a result, RPSS offers a unique design point—a software-only entropy source capable of producing statistically uniform outputs with min-entropy comparable to leading physical and quantum RNGs, while remaining portable across commodity platforms.
From an engineering perspective, RPSS may be viewed as a form of entropy purification achieved through deterministic computational dynamics rather than statistical post-processing, with the observed uniformity emerging from structured aggregation of skewed timing statistics.
This comparison clarifies that RPSS is not intended to replace dedicated hardware RNGs in all settings, but rather to provide a complementary, self-contained alternative where deployability, architectural simplicity, and intrinsic uniformity are prioritized.
@@ IV Conclusion
The empirical validation confirms our theoretical predictions, including the mode, expectation, and variance of the permutation count distributions, which follow the negative binomial model with success probability $p=1/N!$. Table 4 compares the observed and theoretical modes for $N=4$ and m=1m=144, demonstrating close agreement and confirming that the RPSS framework accurately captures both combinatorial and stochastic aspects of the system. The compound nature of the elapsed-time observable, which incorporates real-world system noise, is a powerful source of unpredictability. The modular uniformization process, in which a large number of microstates collapse onto a small set of computational symbols, acts as a cryptographic extractor, eliminating residual correlations and producing outputs that pass stringent statistical tests. Our analysis of entropy convergence as a function of array size ($N$) and repetition parameter ($m$) provides a practical guide for implementing a self-stabilizing entropy source on commodity hardware.